# Graph Anomaly Detection via Adaptive Test-time Representation Learning across Out-of-Distribution Domains

### v1

@@ 6 Ablation Studies
Self-supervised learning during the training phase. To highlight the effectiveness of self-supervision in source training, we compared two scenarios in Table 4: one where the source model was trained using only the supervised loss (AdaGraph-T3-sup-source) and another where it used both supervised and self-supervised losses (AdaGraph-T3). We then evaluated how these different source model training approaches affected the performance of the target model. The results demonstrate that using self-supervision during the training phase is often beneficial. Self-supervision leads to a higher average performance and greater improvement compared to the best baseline (GTrans).
7 Conclusion
In this paper, we have introduced AdaGraph-T3, a novel test-time training framework for graph anomaly detection when the testing data is out-of-distribution and from heterogeneous application domains. Specifically, our approach addresses scenarios where the distributions and feature spaces differ between the source and target datasets using dataset-specific encoders. AdaGraph-T3 combines the advantages of test-time training with multiple innovations focused on graph anomaly detection. For instance, by leveraging the distinctive distribution pattern of node homophily between normal and anomalous nodes, we have presented a tailored SSL task along with a robust attention-based edge-weighting mechanism (NSAW) to enhance the generalization of the learned representations. Moreover, AdaGraph-T3 performs model election using an MMD-based early-stopping criterion. Our experiments have illustrated that AdaGraph-T3 outperforms state-of-the-art graph domain adaptation baselines.
Our work opens several avenues for future investigation. We are interested in applying AdaGraph-T3 to cybersecurity applications, where label scarcity and distribution shifts are a challenge. Moreover, we want to understand how explainers trained in the source domain can improve the interpretability of target anomalies. Finally, we will study theoretical guarantees for cross-domain GAD based on the similarity between source and target datasets.
Appendix A Early stopping criterion
We propose an adaptive early-stopping strategy that monitors distribution shifts between source and target domains using Maximum Mean Discrepancy (MMD). Let $f_{t}\in\mathbb{R}^{n_{t}\times p}$ denote the target embedding, $f_{s}^{n}\in\mathbb{R}^{n_{n}\times p}$ and $f_{s}^{a}\in\mathbb{R}^{n_{a}\times p}$ denote the normal and attack source embeddings respectively, where $n_{t}$, $n_{n}$, $n_{a}$ are the number of samples and $p$ is the common feature dimension. The MMD ratio score at epoch $t$ is computed as:
$\begin{split}\text{MMD}_{n}(t)&=\text{MMD}(f_{t},f_{s}^{n})\\ \end{split}$ (15)
$\text{MMD}_{a}(t)=\text{MMD}(f_{t},f_{s}^{a})$ (16)
$\text{score}(t)=\frac{1}{|\mathcal{V}|}\sum_{i\in\mathcal{V}}\frac{\max(\text{MMD}_{n}^{i},\text{MMD}_{a}^{i})}{\min(\text{MMD}_{n}^{i},\text{MMD}_{a}^{i})}$ (17)
For target attack samples, $\text{MMD}_{n}(t)$ would be larger than $\text{MMD}_{a}(t)$, making the score $\text{MMD}_{n}(t)/\text{MMD}_{a}(t)$, which increases as samples align better with attack source features. Conversely, for target normal samples, the score becomes $\text{MMD}_{a}(t)/\text{MMD}_{n}(t)$, increasing as samples align with normal source features. The training stops at epoch $T$ if:
Appendix B Data description
Amazon [25]: The Amazon dataset consists of product reviews from the Musical Instruments category. Users with more than 80% helpful votes are labeled as benign entities, while those with less than 20% helpful votes are considered fraudulent entities. For each user (represented as a node in the graph), 25 handcrafted features are used as raw node features. The graph structure is defined by the U-P-U (User-Product-User) relation, which connects users who have reviewed at least one common product.
Reddit [16]: The Reddit dataset consists of forum posts from the Reddit platform, focusing on user behavior and content. In this dataset, users who have been banned from the platform are labeled as anomalies. Each post’s textual content has been vectorized to serve as one of the 64 attributes for the corresponding user or post.
Facebook [18]: The Facebook dataset represents a social network structure derived from the Facebook platform. Users establish connections with other users, forming a network of relationships. Fraudulent users are assumed to be anomalies of the network. Each node represents one user with 576 node attributes.
YelpChi [27]: The Yelp spam review dataset comprises hotel and restaurant reviews, categorized as either filtered (spam) or recommended (legitimate) by Yelp. We utilize 32 handcrafted features as raw node features for each review in the Yelp dataset. In the graph the reviews serve as nodes. The graph’s structure is defined by the R-U-R (Review-User-Review) relation, which connects reviews posted by the same user.
YelpHotel [8]: A graph dataset focused on hotel reviews from Yelp, where users and hotels are nodes connected by review edges. Each review contains ratings, text, and detailed metadata about both hotels and reviewers.
YelpRes [8]: A graph dataset containing restaurant reviews from Yelp, where users and restaurants are nodes connected by review edges. Each review has ratings, text, and metadata about both users and restaurants
Appendix C Cross-domain performance
Tables 6 and 7 present the cross-domain and in-domain adaptation performance of AdaGraph-T3. The results are shown using AUROC (%) and AUPRC (%) metrics, respectively. We evaluated every possible combination of four real-world datasets (Amazon, Facebook, YelpChi, and Reddit) as source and target domains.
Source Target
Amazon Facebook YelpChi Reddit
Amazon 77.19 91.04 65.54 62.55
Facebook 86.11 95.11 72.30 60.96
YelpChi 74.02 93.07 73.30 61.29
Reddit 80.56 94.84 70.59 64.78
Table 6: Cross-domain anomaly detection performance of AdaGraph-T3 (AUROC %)
Source Target
Amazon Facebook YelpChi Reddit
Amazon 15.19 31.71 8.83 6.02
Facebook 29.19 29.38 16.67 4.97
YelpChi 15.87 33.89 17.38 5.31
Reddit 19.02 34.76 13.36 6.97
Table 7: Cross-domain anomaly detection performance of AdaGraph-T3 (AUPRC %)
Appendix D Transfer performance with respect to source-target homophily
@@
Appendix E Additional embedding visualizations
@@
In Figure 7, we show 2D embeddings of the learned model for different pairs of source and target (similar to Figure 4).

### v2

@@ A.3 Hyperparameter tuning
A.5 Transfer performance with respect to source-target homophily
@@
A.6 Evaluation using $\theta_{\text{pred}}$ for anomaly scoring
In addition to our ranking-based anomaly scoring, we evaluated an alternative setup where the frozen source-domain classifier $\theta_{\text{pred}}$ is used directly on the adapted target embeddings. Specifically, the predicted class probabilities were used to rank nodes by their likelihood of being anomalous. The same test-time training and early stopping procedures were applied. Performance results using this method are summarized in Table 7, allowing direct comparison with our primary ranking-based approach.
Table 7: Cross-domain node-level anomaly detection results for GADT3 using frozen prediction layer and original GADT3.
Metric (%) Method $\toRDT AMZ\toFB$ $\toAMZ RDT\toFB$ $\toAMZ FB\toRDT$ Avg.
AUROC GADT3 (pred. layer) 59.09 94.80 80.76 94.98 74.39 64.25 78.05
GADT3 61.95 91.03 79.87 94.84 82.57 61.93 78.70
AUPRC GADT3 (pred. layer) 4.48 28.27 29.17 31.98 14.53 6.73 19.19
GADT3 5.19 27.06 19.10 34.76 25.73 6.12 19.99
A.7 Complexity analysis
The per-layer time complexity is $\mathcal{O}(|\mathcal{E}|\cdot d+|\mathcal{V}|\cdot d^{2})$, where $|\mathcal{V}|$ and $|\mathcal{E}|$ are the number of nodes and edges, respectively, and $d$ is the hidden dimension of the projected space Li et al. (2025). Message passing is linear in the number of edges, and our attention mechanism—computed only over each node’s neighbors—adds an extra term in $\mathcal{O}(|\mathcal{E}|)$, without incurring the $\mathcal{O}(|\mathcal{V}|^{2})$ cost of dense similarity computation. Memory complexity is $\mathcal{O}(|\mathcal{V}|\cdot d+|\mathcal{E}|)$ in the standard setting. During test-time adaptation (TTA), only the projection layer for the target domain is updated while the GNN and classifier remain frozen, reducing memory overhead for gradients. Runtime and memory usage with and without NSAW are provided in Table 8.
Table 8: Runtime and memory usage during TTA and additional overhead from NSAW.
Dataset TTA time(s) TTA Memory (MB) NSAW Time (s) NSAW Memory (MB)
AMZ 14.35 2264.17 0.01 0.21
FB 2.34 244.45 0.01 0.30
RDT 14.93 2630.33 0.02 0.14
HTL 10.41 3594.84 0.12 151.29
RES 16.29 3418.60 0.44 17.40
AMZ-all 52.13 3802.78 0.89 6.68
YC-all 50.47 4104.21 1.35 13.16
TF 190.62 7506.49 1.89 18.12
A.8 Additional hyperparameter sensitivity analysis
[…]
$p=32$ 90.1 81.6
$\mathbf{p=40}$ 91.0 82.6
$p=64$ 90.8 81.7
$\lambda_{\text{reg}}=0.05$ 90.2 81.0
$\bm{\lambda_{\text{reg}}=0.1}$ 91.0 82.6
$\lambda_{\text{reg}}=0.2$ 89.8 81.8
$\lambda_{s}=0.0005$ 90.4 81.3
$\bm{\lambda_{s}=0.001}$ 91.0 82.6
$\lambda_{s}=0.005$ 90.5 81.7
A.9 Component Contribution under Homophily and Feature Shift
[…]
Table 10: Component-wise ablation (AUROC %) under different homophily (0.9 vs. 0.3) and feature shift.
Component High Homophily Low Homophily Large Feature Shift Small Feature Shift
[…]
w/o NSAW 61.35 56.27 61.35 81.63
w/o Source 61.54 57.11 61.54 84.76
w/o ClassReg 61.76 55.91 61.76 81.98
GADT3 (full) 61.95 58.30 61.95 82.57
A.10 Proof of Proposition 1
The proof proceeds in two main steps. First, we establish by induction that the gradient dominance property, $\|g_{N}^{(k)}\|>\|g_{A}^{(k)}\|$, holds for all TTT steps. We use this property to show that the margin $\Delta_{t}^{(k)}$ is increasing.
We will prove by induction that $\|g_{N}^{(k)}\|>\|g_{A}^{(k)}\|$ for all $k\geq 0$. For the base case $k=0$, we need to show that $\|g_{N}^{(0)}\|>\|g_{A}^{(0)}\|$. By Assumption 2, we have $\|g_{C}^{(0)}-g_{C}(\theta_{s}^{*})\|\leq L\|\theta_{t}^{(0)}-\theta_{s}^{*}\|$. By the triangle inequality,
$\|g_{N}^{(0)}\|-\|g_{A}^{(0)}\|\geq\|g_{N}(\theta_{s}^{*})\|-\|g_{A}(\theta_{s}^{*})\|-\|g_{N}^{(0)}-g_{N}(\theta_{s}^{*})\|-\|g_{A}^{(0)}-g_{A}(\theta_{s}^{*})\|.$
From Assumption 1, $\|g_{N}(\theta_{s}^{*})\|-\|g_{A}(\theta_{s}^{*})\|=\varepsilon_{0}>0$. By choosing a sufficiently small $\delta$ such that $\|\theta_{t}^{(0)}-\theta_{s}^{*}\|<\delta$ (Assumption 3), we can make $\|g_{N}^{(0)}-g_{N}(\theta_{s}^{*})\|<L\delta$ and $\|g_{A}^{(0)}-g_{A}(\theta_{s}^{*})\|<L\delta$. Thus, by choosing $\delta$ small enough, we ensure $\|g_{N}^{(0)}\|-\|g_{A}^{(0)}\|>0$. This establishes the base case.
For the inductive step, assume that for some $k\geq 0$, the hypothesis holds: $\|g_{N}^{(k)}\|>\|g_{A}^{(k)}\|$. We must show that this implies $\|g_{N}^{(k+1)}\|>\|g_{A}^{(k+1)}\|$. By definition, the global TTT gradient at step $k$ is
$g_{V}(\theta^{(k)}_{t})=\nabla_{\theta}\left.\mathbb{E}_{v\in\mathcal{V}}[s_{t}(v;\theta)]\right|_{\theta=\theta_{t}^{(k)}}=r_{N}g_{N}(\theta_{t}^{(k)})+r_{A}g_{A}(\theta_{t}^{(k)}),$
where $r_{N}=|\mathcal{N}|/|\mathcal{V}|$ and $r_{A}=1-r_{N}$. By the nature of graph anomaly detection, the number of normal nodes is typically much larger than the number of anomalous nodes, so $r_{N}>r_{A}$. This implies that $(r_{N}-r_{A})>0$. Using the shorthand notations introduced earlier, this can be written as $g_{V}^{(k)}=r_{N}g_{N}^{(k)}+r_{A}g_{A}^{(k)}$.
The TTT update is $\theta_{t}^{(k+1)}=\theta_{t}^{(k)}+\eta g_{V}^{(k)}$. By Assumption 2, we have $\|g_{C}^{(k+1)}-g_{C}^{(k)}\|\leq L\|\theta_{t}^{(k+1)}-\theta_{t}^{(k)}\|=L\eta\|g_{V}^{(k)}\|$. Using the triangle inequality, we can bound the norms of the gradients at step $k+1$:
$\|g_{N}^{(k+1)}\|$ $\geq\|g_{N}^{(k)}\|-\|g_{N}^{(k+1)}-g_{N}^{(k)}\|\geq\|g_{N}^{(k)}\|-L\eta\|g_{V}^{(k)}\|$
$\|g_{A}^{(k+1)}\|$ $\leq\|g_{A}^{(k)}\|+\|g_{A}^{(k+1)}-g_{A}^{(k)}\|\leq\|g_{A}^{(k)}\|+L\eta\|g_{V}^{(k)}\|$
Subtracting the second inequality from the first gives
$\|g_{N}^{(k+1)}\|-\|g_{A}^{(k+1)}\|\geq\left(\|g_{N}^{(k)}\|-\|g_{A}^{(k)}\|\right)-2L\eta\|g_{V}^{(k)}\|.$
By the inductive hypothesis, $\|g_{N}^{(k)}\|>\|g_{A}^{(k)}\|$. Given the boundedness gradient assumption, we can choose a small enough $\eta$ such that the above term is positive, ensuring that the gradient dominance is preserved. This concludes the induction.
Now we can use this property to prove the proposition. A first-order Taylor expansion of $\Delta_{t}$ at $\theta_{t}^{(k)}$ gives
$\Delta_{t}^{(k+1)}-\Delta_{t}^{(k)}=\eta\,\langle g_{N}^{(k)}-g_{A}^{(k)},g_{V}^{(k)}\rangle+O(\eta^{2}).$
Substituting this into the inner product, we get
$\langle g_{N}^{(k)}-g_{A}^{(k)},r_{N}g_{N}^{(k)}+r_{A}g_{A}^{(k)}\rangle$ $=r_{N}\|g_{N}^{(k)}\|^{2}-r_{A}\|g_{A}^{(k)}\|^{2}+(r_{A}-r_{N})\langle g_{N}^{(k)},g_{A}^{(k)}\rangle$
$\geq r_{N}\|g_{N}^{(k)}\|^{2}-r_{A}\|g_{A}^{(k)}\|^{2}-(r_{N}-r_{A})\|g_{N}^{(k)}\|\|g_{A}^{(k)}\|$
$>r_{N}\|g_{N}^{(k)}\|^{2}-r_{A}\|g_{N}^{(k)}\|^{2}-(r_{N}-r_{A})\|g_{N}^{(k)}\|^{2}$
$=0$
For the first inequality, we apply the Cauchy-Schwarz inequality, which gives $|\langle g_{N}^{(k)},g_{A}^{(k)}\rangle|\leq\|g_{N}^{(k)}\|\|g_{A}^{(k)}\|$. Moreover, from the inductive proof, we have $\|g_{N}^{(k)}\|>\|g_{A}^{(k)}\|$ hence the second inequality holds. Given that all terms in the final expansion are positive, the inner product term is guaranteed to be positive. For a sufficiently small $\eta$, the quadratic term $O(\eta^{2})$ is negligible. The sign of $\Delta_{t}^{(k+1)}-\Delta_{t}^{(k)}$ is determined by the sign of the inner product term, which is positive. Thus, we have $\Delta_{t}^{(k+1)}>\Delta_{t}^{(k)}$ which concludes the proof.
A.11 Homophily score patterns across datasets
Figures 8 and 9 show homophily score distributions for additional datasets.
Figure 8: Homophily score distributions across domains (Reddit and Facebook)
Figure 9: Homophily score distributions across domains (YelpRes and YelpHotel)
A.12 Additional embedding visualizations
@@
In Figure 10, we show 2D embeddings from our model for different sources and targets (similar to Figure 4).